OpenStudy (lonababy):

pleaseeeeeeee help me.Brady has a collection of dimes and quarters totaling $32.85. she has 171 coins in all. how many of each coin does she have?

OpenStudy (texaschic101):

d + q = 171 ----> d = 171 - q .10d + .25q = 32.85 now we will use substitution and plug in 171 - q in for d in the 2nd equation .10(171 - q) + .25q = 32.85 -- distribute through the parenthesis 17.1 - .10q + .25q = 32.85 --- subtract 17.1 from both sides -.10q + .25q = 32.85 - 17.1 -- combine terms 0.15q = 15.75 -- divide both sides by 0.15 q = 15.75/0.15 q = 105 d + q = 171 d + 105 = 171 d = 171 - 105 d = 66 check.. .10d + .25q = 32.85 .10(66) + .25(105) = 32.85 6.60 + 26.25 = 32.85 32.85 = 32.85 (correct) there are 105 quarters (q) and 66 dimes (d)

OpenStudy (lonababy):

oh wow thanks I understand it now